OpenStudy (anonymous):

1. Call Wireless phone company charges $0.50 for the first call you make and $0.10 for each additional minute. The cost of a phone call can be expressed by the formula C = 0.50 + 0.10(m – 1), where C is the total cost in dollars and m is the number of minutes that you talk. If the bill for your phone call was $1.90, how long did you talk? A. 12 minutes B. 14 minutes C. 15 minutes D. 16 minutes 2. You spend $77 shopping. You buy some shirts at $12 each, and spend $5 on food. How many shirts did you buy? A. 5x + 12 = 77 B. x(12 + 5) = 77 C. 12x + 5 = 77 D. 12 X 5+x=77

OpenStudy (igreen):

For #1 we can plug in 1.90 into the equation: 1.90 = 0.50 + 0.10(m - 1) Distribute 0.10 into the parenthesis: 1.90 = 0.50 + 0.10m - 0.10 Simplify: 1.90 = 0.40 + 0.10m Subtract 0.40 to both sides, what's 1.90 - 0.40? @courtney-ray

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

and and with number 2?

OpenStudy (igreen):

That's not your answer yet..we get: 1.5 = 0.10m Now we divide 0.10 to both sides, what's 1.5 / 0.10?

OpenStudy (igreen):

@courtney-ray

OpenStudy (anonymous):

it was 15

OpenStudy (anonymous):

all i had to do was take out the decimal

OpenStudy (igreen):

Yep, 15 is your answer for #1.

OpenStudy (igreen):

Now for #2 '5' will be by itself since it's a 1-time thing, and 12 will have a variable because that's the price PER shirt, and the variable will represent the # of shirts you buy. Then we will have = 77 because that's the amount you spent. So which one looks most like this? @courtney-ray

OpenStudy (anonymous):

its B

OpenStudy (igreen):

No, B will put a variable on 5 too, but we don't want a variable on 5, we just want it to be alone.

OpenStudy (anonymous):

well i entered B and it said it was right

OpenStudy (igreen):

Oh..well actually it should be C..

OpenStudy (igreen):

12x + 5 = 77 Subtract 5 to both sides: 12x = 72 Divide 12 to both sides: x = 6 So you bought 6 shirts.. Check: 12(6) + 5 = 77 72 + 5 = 77 77 = 77

OpenStudy (igreen):

x(12 + 5) = 77 Distribute 'x' into the parenthesis: 12x + 5x = 77 Add 12x + 5x: 17x = 77 Divide 17 to both sides: x = 4.52941176 You can't buy 4.52941176 shirts..so B is obviously wrong..